# Samba: Simple Hybrid State Space Models for Efficient Unlimited Context Language Modeling

### v1

@@ Abstract
Efficiently modeling sequences with infinite context length has been a long-standing problem. Past works suffer from either the quadratic computation complexity or the limited extrapolation ability on length generalization. In this work, we present Samba, a simple hybrid architecture that layer-wise combines Mamba, a selective State Space Model (SSM), with Sliding Window Attention (SWA). Samba selectively compresses a given sequence into recurrent hidden states while still maintaining the ability to precisely recall memories with the attention mechanism. We scale Samba up to 3.8B parameters with 3.2T training tokens and show that Samba substantially outperforms the state-of-the-art models based on pure attention or SSMs on a wide range of benchmarks. When trained on 4K length sequences, Samba can be efficiently extrapolated to 256K context length with perfect memory recall and show improved token predictions up to 1M context length. As a linear-time sequence model, Samba enjoys a 3.73×3.73$\times$3.73 × higher throughput compared to Transformers with grouped-query attention when processing user prompts of 128K length, and 3.64×3.64$\times$3.64 × speedup when generating 64K tokens with unlimited streaming. A sample implementation of Samba is publicly available in https://github.com/microsoft/Samba.
@@ 1 Introduction
(a) Perplexity on the test set of Proof-Pile
(b) Decoding throughput with a batch size of 16
Figure 1: Samba shows improved prediction up to 1M tokens in the Proof-Pile test set while achieving a 3.64$\times$ faster decoding throughput than the Llama-3 architecture [Met24] (a state-of-the-art Transformer [VSP+17] with Grouped-Query Attention [ALTdJ+23]) on 64K generation length. We also include an SE-Llama-3 1.6B baseline which applies the SelfExtend [JHY+24] approach for zero-shot length extrapolation. Throughput measured on a single A100 80GB GPU. All models are trained on the Phi-2 [LBE+23] dataset with 4K sequence length.
@@
Attention-based models [VSP+17, BCB14] have dominated the neural architectures of Large Language Models (LLMs) [RWC+19, BMR+20, Ope23, BCE+23] due to their ability to capture complex long-term dependencies and the efficient parallelization for large-scale training [DFE+22]. Recently, State Space Models (SSMs) [GGR21, SWL23, GGGR22, GD23] have emerged as a promising alternative, offering linear computation complexity and the potential for better extrapolation to longer sequences than seen during training. Specifically, Mamba[GD23], a variant of SSMs equipped with selective state spaces, has demonstrated notable promise through strong empirical performance and efficient hardware-aware implementation. Recent work also shows that transformers have poorer modeling capacities than input-dependent SSMs in state tracking problems [MPS24]. However, SSMs struggle with memory recall due to their Markovian nature [AET+23], and experimental results on information retrieval-related tasks [FDS+23, WDL24, AEZ+24], have further shown that SSMs are not as competitive as their attention-based counterparts.
@@
Previous works [ZLJ+22, FDS+23, MZK+23, RLW+23] have explored different approaches to hybridize SSMs and the attention mechanism, but none of them achieve unlimited-length extrapolation with linear-time complexity. The existing length generalization techniques [HWX+23, XTC+23, JHY+24] developed for the attention mechanism suffer from quadratic computation complexity or limited context extrapolation ability. In this paper, we introduce Samba, a simple neural architecture that harmonizes the strengths of both the SSM and the attention-based models, while achieving an unlimited sequence length extrapolation with linear time complexity. Samba combines SSMs with attention through layer-wise interleaving Mamba [GD23], SwiGLU [Sha20], and Sliding Window Attention (SWA) [BPC20]. Mamba layers capture the time-dependent semantics and provide a backbone for efficient decoding, while SWA fills in the gap modeling complex, non-Markovian dependencies.
@@
We scale Samba with 421M, 1.3B, 1.7B and up to 3.8B parameters. In particular, the largest 3.8B base model pre-trained with 3.2T tokens achieves a 71.2 score for MMLU [HBB+21], 54.9 for HumanEval [CTJ+21], and 69.6 for GSM8K [CKB+21], substantially outperforming strong open source language models up to 8B parameters, as detailed in Table 1. Despite being pre-trained in the 4K sequence length, Samba can be extrapolated to 1M length in zero shot with improved perplexity on Proof-Pile [ZAP22] while still maintaining the linear decoding time complexity with unlimited token streaming, as shown in Figure 1. We show that when instruction-tuned in a 4K context length with only 500 steps, Samba can be extrapolated to a 256K context length with perfect memory recall in Passkey Retrieval [MJ23]. In contrast, the fine-tuned SWA-based model simply cannot recall memories beyond 4K length. We further demonstrate that the instruction-tuned Samba 3.8B model can achieve significantly better performance than the SWA-based models on downstream long-context summarization tasks, while still keeping its impressive performance on the short-context benchmarks. Finally, we conduct rigorous and comprehensive analyzes and ablation studies, encompassing up to 1.7 billion parameters, to validate the architectural design of Samba. These meticulous investigations not only justify our architectural designs but also elucidate the potential mechanisms underpinning the remarkable effectiveness of this simple hybrid approach.
@@ 2 Methodology
We explore different hybridization strategies consisting of the layers of Mamba, Sliding Window Attention (SWA), and Multi-Layer Perceptron [Sha20, DFAG16]. We conceptualize the functionality of Mamba as the capture of recurrent sequence structures, SWA as the precise retrieval of memory, and MLP as the recall of factual knowledge. We also explore other linear recurrent layers including Multi-Scale Retention [SDH+23] and GLA [YWS+23] as potential substitutions for Mamba in Section 3.2. Our goal of hybridization is to harmonize between these distinct functioning blocks and find an efficient architecture for language modeling with unlimited-length extrapolation ability.
@@ 2.1 Architecture
Figure 2: From left to right: Samba, Mamba-SWA-MLP, Mamba-MLP, and Mamba. The illustrations depict the layer-wise integration of Mamba with various configurations of Multi-Layer Perceptrons (MLPs) and Sliding Window Attention (SWA). We assume the total number of intermediate layers to be $N$, and omit the embedding layers and output projections for simplicity. Pre-Norm [XYH+20, ZS19] and skip connections [HZRS16] are applied for each of the intermediate layers.
@@ 2.1.1 Mamba Layer
Mamba [GD23] is a recently proposed SSM-based model with selective state spaces. It enables input-dependent gating to both the recurrent states and the input representation for a soft selection of the input sequence elements. Given an input sequence representation $\mathbf{X}\in\mathbb{R}^{n\times d_{m}}$, where $n$ is the length of the sequence and $d_{m}$ is the hidden size, Mamba first expands the inputs to a higher dimension $d_{e}$, i.e.,
@@
where $\mathbf{W}_{\text{in}}\in\mathbb{R}^{d_{m}\times d_{e}}$ is a learnable projection matrix. Then a Short Convolution (SC) [PMN+23] operator is applied to smooth the input signal,
@@
where $\mathbf{W}_{\text{conv}}\in\mathbb{R}^{k\times d_{e}}$ and the kernel size $k$ is set to 4 for hardware-aware efficiency. The Depthwise Convolution [HQW+19] is applied over the sequence dimension followed by a SiLU [EUD17] activation function. The selective gate is then calculated through a low-rank projection followed by Softplus [ZYL+15],
@@
where $\mathbf{Z}_{0}=\mathbf{0}$, $\odot$ means the point-wise product, $\otimes$ means the outer product and $\exp$ means the point-wise natural exponential function. $\mathbf{D}\in\mathbb{R}^{d_{e}}$ is a learnable vector initialized as $D_{i}=1$ and $\mathbf{A}\in\mathbb{R}^{d_{e}\times d_{s}}$ is a learnable matrix initialized as $A_{ij}=\log(j),1\leq j\leq d_{s}$, following the S4D-Real [GGGR22] initialization. In practice, Mamba implements a hardware-aware parallel scan algorithm for efficient parallelizable training. The final output is obtained through a gating mechanism similar to Gated Linear Unit [Sha20, DFAG16],
@@ 2.1.2 Sliding Window Attention (SWA) Layer
The Sliding Window Attention [BPC20] layer is designed to address the limitations of the Mamba layer in capturing non-Markovian dependencies in sequences. Our SWA layer operates on a window size $w=2048$ that slides over the input sequence, ensuring that the computational complexity remains linear with respect to the sequence length. The RoPE [SLP+21] relative positions are applied within the sliding window. By directly accessing the contents in the context window through attention, the SWA layer can retrieve high-definition signals from the middle to short-term history that cannot be clearly captured by the recurrent states of Mamba. We use FlashAttention 2 [Dao23] for the efficient implementation of self-attention throughout this work. We also choose the 2048 sliding window size for efficiency consideration; FlashAttention 2 has the same training speed as Mamba’s selective parallel scan at the sequence length of 2048 based on the measurements in [GD23].
@@ 2.1.3 Multi-Layer Perceptron (MLP) Layer
The MLP layers in Samba serve as the architecture’s primary mechanism for nonlinear transformation and recall of factual knowledge [DDH+22]. We use SwiGLU [Sha20] for all the models trained in this paper and denote its intermediate hidden size as $d_{p}$. As shown in Figure 2, Samba applies separate MLPs for different types of information captured by Mamba and the SWA layers.
@@ 3 Experiments and Results
We pre-train four Samba models with different parameter sizes, 421M, 1.3B, 1.7B and 3.8B, to investigate its performance across different scales. The details of the hyperparameters for the training and architecture designs are shown in Table 10 of Appendix A. We also train other hybrid architectures as mentioned in Section 2.1, including the baseline Mamba, Llama-3, and Mistral architecture on a scale of around 1.7B, with detailed hyperparameters in Table 9 of Appendix A. We do comprehensive downstream evaluations on a wide range of benchmarks, focusing on four main capabilities of the models: commonsense reasoning (ARC [CCE+18], PIQA [BZB+20], WinoGrande [SBBC21], SIQA [SRC+19]), language understanding (HellaSwag [ZHB+19], BoolQ [CLC+19], OpenbookQA [MCKS18], SQuAD [RZLL16], MMLU [HBB+21]), truthfulness (TruthfulQA [LHE22]) and math and coding (GSM8K [CKB+21], MBPP [AON+21], HumanEval [CTJ+21]).
@@ 3.1 Language Modeling on Textbook Quality Data
We first present results from our largest 3.8B Samba model, trained on the same data set used by Phi3 [AJA+24] with 3.2T tokens. We follow the same multi-phase pretraining strategy as Phi3-mini for a fair comparison. We also report the performance of the Transformer++ (TFM++ in Table 1) model, which uses the same architecture and training recipe as Phi3-mini, for a fair comparison. In Table 1, we conduct comprehensive evaluations on a diverse subset of the benchmarks to assess Samba’s performance across all the domains mentioned above to ensure a thorough examination of the model’s capabilities. The details of the generation configurations are included in Appendix A.
Table 1: Downstream performance comparison of Samba 3.8B with other pretrained base language models without instruction tuning. ARC-C and HellaSwag are measured with character-normalized accuracy. MMLU and GSM8K are measured in 5-shot, while others are in zero-shot. We report the MC2 score for TruthfulQA, maj@1 for GSM8K, and pass@1 for HumanEval. Measured by ours.
Model Size Tokens MMLU Hella- ARC- Wino- Truth. GSM Hum. Avg.
Swag C Gran. QA 8K Eval
Llama 2 6.7B 2T 45.3 77.2 45.9 69.2 38.8 14.6 12.8 43.4
13B 2T 54.8 80.7 49.4 72.8 37.4 28.7 18.3 48.9
Mistral 7.2B - 60.1 81.3 55.5 75.3 42.2 35.4 30.5 53.6
Mamba 2.8B 600B 26.2 71.0 41.7 65.9 34.4 3.6 7.3 35.7
Gemma 2.5B 3T 42.3 71.4 42.1 65.4 33.1 17.7 22.0 42.0
8.5B 6T 64.3 81.2 53.2 72.3 44.8 46.4 32.3 56.4
R-Gemma 2.7B 2T 38.4 71.0 42.3 67.8 35.1 13.4 21.3 41.3
Llama 3 8.0B 15T+ 66.6 79.2 53.2 72.6 43.9 45.8 28.7 55.8
TFM++ 3.8B 3.2T 67.2 76.6 53.8 72.6 47.3 51.5 51.8 60.1
Samba 3.8B 3.2T 71.2 77.4 55.7 77.1 43.4 69.6 54.9 64.2
We compare with several strong baselines, including Llama 2 [TMS+23], Mistral [JSM+23], Mamba [GD23], Gemma [Tea24], Recurrent-Gemma (R-Gemma) [BDS+24], Llama 3 [Met24] and TFM++. As shown in Table 1, Samba achieves the highest average score on all benchmarks, demonstrating its superior performance in handling various language comprehension tasks. Notably, Samba excels in the GSM8K benchmark, achieving an absolute 18.1% higher accuracy than TFM++ trained on the same dataset. This shows the surprising complementary effect of combining SSM with the attention mechanism. We conjecture that when combined with attention, Mamba, as an input-dependent SSM, can focus more on performing the arithmetic operation through its recurrent states than on doing the retrieval operation which can be easily learned by the sliding window attention.
@@
Benchmark Llama-3 Mistral Mamba Mamba-SWA- Mamba- Samba
1.6B 1.6B 1.8B MLP 1.6B MLP 1.9B 1.7B
@@
To examine the different hybridization strategies mentioned in Section 2.1, we train 6 models with around 1.7B parameters on the Phi2 [LBE+23] dataset with 230B tokens and evaluate them in the full suite of 15 downstream benchmarks to have a holistic assessment of hybrid and purebred architectures. As shown in Table 2, Samba demonstrates superior performance on a diverse set of tasks, including commonsense reasoning (ARC-Challenge), language understanding (MMLU, SQuAD), TruthfulQA and code generation (HumanEval, MBPP). It outperforms both the pure attention-based and SSM-based models in most tasks and achieves the best average performance. We can observe that replacing Mamba blocks with MLPs does not harm commonsense reasoning ability, but its performance on language understanding and complex reasoning ability, such as coding and mathematical reasoning, degenerates significantly. We can also see that pure Mamba models fall short on retrieval intensive tasks such as SQuAD due to their lack of precise memory retrieval ability. The best results are achieved through the combination of the attention and Mamba modules, as shown with our Samba architecture. We can also notice that Mamba-SWA-MLP has significantly better performance on GSM8K, potentially resulting from a closer collaboration between the Mamba and the SWA layers. The distinct downstream performances of different hybridization strategies pose interesting future work for developing task-adaptive dynamic architectures.
3.2 Exploration on Attention and Linear Recurrence
Since SSMs belong to a broader realm of linear recurrent models [OSG+23, QYZ23, YWS+23, Kat23, QYS+24], there exist multiple alternatives other than Mamba when combing attention-based layers with recurrent neural networks. In addition to Mamba and Samba, we investigate the comparative analysis of the following architectures:
  • Llama-2 [TMS+23] is an attention-based Transformer architecture that utilizes full self-attention across the entire sequence.
  • Llama-2-SWA is an attention-based architecture that replaces all full attention layers in Llama-2 with sliding window attention.
  • Sliding RetNet replaces Mamba layers in the Samba architecture with Multi-Scale Retention  [SDH+23] layers. RetNet is a linear attention model with fixed and input-independent decay applying to the recurrent hidden states.
  • Sliding GLA replaces Mamba layers in the Samba architecture with Gated Linear Attention (GLA) [YWS+23]. GLA is a more expressive variant of linear attention with input-dependent gating.

### v2

@@ Abstract
Efficiently modeling sequences with infinite context length has long been a challenging problem. Previous approaches have either suffered from quadratic computational complexity or limited extrapolation ability in length generalization. In this work, we present Samba, a simple hybrid architecture that layer-wise combines Mamba, a selective State Space Model (SSM), with Sliding Window Attention (SWA). Samba selectively compresses a given sequence into recurrent hidden states while still maintaining the ability to precisely recall recent memories with the attention mechanism. We scale Samba up to 3.8B parameters with 3.2T training tokens and demonstrate that it significantly outperforms state-of-the-art models across a variety of benchmarks. Pretrained on sequences of 4K length, Samba shows improved perplexity in context lengths of up to 1M in zero-shot. When finetuned on 4K-length sequences, Samba efficiently extrapolates to a 256K context length with perfect memory recall on the Passkey Retrieval task, and exhibits superior retrieval extrapolation on the challenging Phonebook task compared to full-attention models. As a linear-time sequence model, Samba achieves a 3.73×3.73$\times$3.73 × higher throughput compared to Transformers with grouped-query attention for user prompts of 128K length, and a 3.64×3.64$\times$3.64 × speedup when generating 64K tokens with unlimited streaming. Our code for training on open source data is publicly available at https://github.com/microsoft/Samba.
@@ 1 Introduction
Attention-based models (Vaswani et al., 2017; Bahdanau et al., 2014) have dominated the neural architectures of Large Language Models (LLMs) (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023; Bubeck et al., 2023) due to their ability to capture complex long-term dependencies and the efficient parallelization for large-scale training (Dao et al., 2022a). Recently, State Space Models (SSMs) (Gu et al., 2021; Smith et al., 2023; Gu et al., 2022; Gu & Dao, 2023) have emerged as a promising alternative, offering linear computation complexity and the potential for better extrapolation to longer sequences than seen during training. Specifically, Mamba (Gu & Dao, 2023), a variant of SSMs equipped with selective state spaces, has demonstrated notable promise through strong empirical performance and efficient hardware-aware implementation. Recent work also shows that transformers have poorer modeling capacities than input-dependent SSMs in state tracking problems (Merrill et al., 2024). However, SSMs struggle with memory recall due to their recurrent nature (Arora et al., 2023), and experimental results on information retrieval-related tasks (Fu et al., 2023; Wen et al., 2024; Arora et al., 2024), have further shown that SSMs are not as competitive as their attention-based counterparts.
@@
Previous works (Zuo et al., 2022; Fu et al., 2023; Ma et al., 2023; Ren et al., 2023) have explored various approaches to hybridize SSMs with the attention mechanism, but none have demonstrated significantly better language modeling performance compared to state-of-the-art Transformer architectures. Existing length extrapolation techniques (Han et al., 2023; Xiao et al., 2023; Jin et al., 2024) designed for attention mechanisms are constrained by quadratic computational complexity or insufficient context extrapolation performance, particularly when evaluated under perplexity metrics. In this paper, we introduce Samba, a simple neural architecture that harmonizes the strengths of both the SSM and the attention-based models, while achieving a potentially infinite length extrapolation with linear time complexity. Samba combines SSMs with attention through layer-wise interleaving Mamba (Gu & Dao, 2023), SwiGLU (Shazeer, 2020), and Sliding Window Attention (SWA) (Beltagy et al., 2020). Mamba layers capture the time-dependent semantics and provide a backbone for efficient decoding, while SWA fills in the gap modeling complex, non-recurrent dependencies. A detailed discussion of related work is included in Appendix A.
@@
We scale Samba with 421M, 1.3B, 1.7B and up to 3.8B parameters with 3.2T tokens. In particular, the largest 3.8B post-trained model achieves a 47.9 score for MMLU-Pro (Hendrycks et al., 2021), 70.1 for HumanEval (Chen et al., 2021), and 86.4 for GSM8K (Cobbe et al., 2021), substantially outperforming strong open source language models up to 8B parameters, as detailed in Table 8. Despite being pre-trained in the 4K sequence length, Samba can be extrapolated to 1M length in zero shot with improved perplexity on Proof-Pile (Zhangir Azerbayev & Piotrowski, 2022), achieving a 256$\times$ extrapolation ratio, while still maintaining the linear decoding time complexity with unlimited token streaming, as shown in Figure 2. We show that when instruction-tuned in a 4K context length with only 500 steps, Samba can be extrapolated to a 256K context length with perfect memory recall in Passkey Retrieval (Mohtashami & Jaggi, 2023). In contrast, the fine-tuned SWA-based model simply cannot recall memories beyond 4K length. We further demonstrate that the instruction-tuned Samba 3.8B model can achieve significantly better performance than the SWA-based models on downstream long-context summarization tasks, while still keeping its impressive performance on the short-context benchmarks. In a more challenging multiple key-value retrieval task, Phonebook (Jelassi et al., 2024), we demonstrate that instruction fine-tuning enables Samba to bridge the retrieval performance gap with full-attention models, while exhibiting significantly better extrapolation ability when retrieving phone numbers beyond the training context length. Finally, we perform extensive analyzes and ablation studies across model sizes up to 1.7B parameters to validate the architectural design of Samba. We also offer potential explanations for the effectiveness of our simple hybrid approach through the lens of attention/selection entropy. To the best of our knowledge, Samba is the first hybrid model showing that linear complexity models can be substantially better than state-of-the-art Transformer models on short-context tasks at large scale, while still being able to extrapolate to extremely long sequences under the perplexity metric.
@@ 2 Methodology
We explore different hybridization strategies consisting of the layers of Mamba, Sliding Window Attention (SWA), and Multi-Layer Perceptron (Shazeer, 2020; Dauphin et al., 2016). We conceptualize the functionality of Mamba as the capture of recurrent sequence structures, SWA as the precise retrieval of memory, and MLP as the recall of factual knowledge. We also explore other linear recurrent layers including Multi-Scale Retention (Sun et al., 2023) and GLA (Yang et al., 2023) as potential substitutions for Mamba in Section 3.2. Our goal of hybridization is to harmonize between these distinct functioning blocks and find an efficient architecture for language modeling with unlimited length extrapolation ability.
Figure 1: From left to right: Samba, Mamba-SWA-MLP, Mamba-MLP, and Mamba. The illustrations depict the layer-wise integration of Mamba with various configurations of Multi-Layer Perceptrons (MLPs) and Sliding Window Attention (SWA). We assume the total number of intermediate layers to be $N$, and omit the embedding layers and output projections for simplicity. Pre-Norm (Xiong et al., 2020; Zhang & Sennrich, 2019) and skip connections (He et al., 2016) are applied for each of the intermediate layers.
@@ 2.1.1 Mamba Layer
Mamba (Gu & Dao, 2023) is a recently proposed SSM-based model with selective state spaces. It enables input-dependent gating to both the recurrent states and the input representation for a soft selection of the input sequence elements. Given an input sequence representation $\mathbf{X}\in\mathbb{R}^{n\times d_{m}}$, where $n$ is the length of the sequence and $d_{m}$ is the hidden size, Mamba first expands the inputs to a higher dimension $d_{e}$, i.e.,
@@
where $\mathbf{W}_{\text{in}}\in\mathbb{R}^{d_{m}\times d_{e}}$ is a learnable projection matrix. Then a Short Convolution (SC) (Poli et al., 2023) operator is applied to smooth the input signal,
@@
where $\mathbf{W}_{\text{conv}}\in\mathbb{R}^{k\times d_{e}}$ and the kernel size $k$ is set to 4 for hardware-aware efficiency. The Depthwise Convolution (He et al., 2019) is applied over the sequence dimension followed by a SiLU (Elfwing et al., 2017) activation function. The selective gate is then calculated through a low-rank projection followed by Softplus (Zheng et al., 2015),
@@
where $\mathbf{Z}_{0}=\mathbf{0}$, $\odot$ means the point-wise product, $\otimes$ means the outer product and $\exp$ means the point-wise natural exponential function. $\mathbf{D}\in\mathbb{R}^{d_{e}}$ is a learnable vector initialized as $D_{i}=1$ and $\mathbf{A}\in\mathbb{R}^{d_{e}\times d_{s}}$ is a learnable matrix initialized as $A_{ij}=\log(j),1\leq j\leq d_{s}$, following the S4D-Real (Gu et al., 2022) initialization. In practice, Mamba implements a hardware-aware parallel scan algorithm for efficient parallelizable training. The final output is obtained through a gating mechanism similar to Gated Linear Unit (Shazeer, 2020; Dauphin et al., 2016),
@@ 2.1.2 Sliding Window Attention (SWA) Layer
We include Sliding Window Attention (Beltagy et al., 2020) layers to address the limitations of Mamba layers in capturing non-recurrent dependencies in sequences. Our SWA layer operates on a window size $w=2048$ that slides over the input sequence, ensuring that the computational complexity remains linear with respect to the sequence length. RoPE (Su et al., 2021) is applied within the sliding window, with a base frequency of 10,000. By directly accessing the contents in the context window through attention, the SWA layer can retrieve high-definition signals from the middle to short-term history that cannot be clearly captured by the recurrent states of Mamba. We use FlashAttention 2 (Dao, 2023) for the efficient implementation of self-attention throughout this work. We also choose the 2048 sliding window size for efficiency consideration; FlashAttention 2 has the same training speed as Mamba’s selective parallel scan at the sequence length of 2048 based on the measurements in (Gu & Dao, 2023).
@@ 2.1.3 Multi-Layer Perceptron (MLP) Layer
The MLP layers in Samba serve as the architecture’s primary mechanism for nonlinear transformation and recall of factual knowledge (Dai et al., 2022). We use SwiGLU (Shazeer, 2020) for all the models trained in this paper and denote its intermediate hidden size as $d_{p}$. As shown in Figure 1, Samba applies separate MLPs for different types of information captured by Mamba and the SWA layers.
@@ 3 Experiments and Results
We pre-train four Samba models with different parameter sizes, 421M, 1.3B, 1.7B and 3.8B, to investigate its performance across different scales. The details of the hyperparameters for the training and architecture designs are shown in Table 12 of Appendix G. We also train other hybrid architectures as mentioned in Section 2.1, including the baseline Mamba (Gu & Dao, 2023), Llama-3 (MetaAI, 2024; Dubey et al., 2024), and Mistral (Jiang et al., 2023) architecture on a scale of around 1.7B, with detailed hyperparameters in Table 11 of Appendix G. We do comprehensive downstream evaluations on a wide range of benchmarks, focusing on four main capabilities of the models: commonsense reasoning (ARC (Clark et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), SIQA (Sap et al., 2019)), language understanding (HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), OpenbookQA (Mihaylov et al., 2018), SQuAD (Rajpurkar et al., 2016), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024), GPQA(Rein et al., 2023)), truthfulness (TruthfulQA (Lin et al., 2022)) and math and coding (GSM8K (Cobbe et al., 2021), MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021)).
Table 1: Downstream performance comparison between Samba-3.8B-IT (preview) and Phi-3-mini-4K on both long-context and short-context tasks. We report 5-shot accuracy (averaged by category) for MMLU, 8-shot CoT (Wei et al., 2022) for GSM8K, 0-shot pass@1 for HumanEval, ROUGE-L for both GovReport and SQuALITY. † Results from the Phi-3 technical report (Abdin et al., 2024).
Model MMLU GSM8K HumanEval GovReport SQuALITY
Phi-3-mini-4K-instruct † 68.8 82.5 58.5 14.4 21.6
Samba-3.8B-IT (preview) 71.9 87.6 62.8 18.9 21.2
@@ 3.1 Language Modeling on Textbook Quality Data
We first present results from our largest 3.8B Samba model, trained on the same data set used by Phi3 (Abdin et al., 2024) with 3.2T tokens. We follow the same multiphase pretraining strategy as Phi3-mini, and apply both the original Phi-3-mini post-training recipe and the Phi3-mini-June-2024 recipe to produce our instruction-tuned Samba 3.8B models, i.e., Samba-3.8B-IT (preview) and Samba-3.8B (June) respectively. We report comprehensive benchmark results of the Samba 3.8B base model and Samba-3.8B (June) in Appendix B. As shown in Table 1, we evaluate the downstream performance of Samba-3.8B-IT (preview) on both long-context summarization tasks (GovReport (Huang et al., 2021), SQuALITY (Wang et al., 2022)) and major short-context benchmarks (MMLU, GSM8K, HumanEval). We can see that Samba has substantially better performance than Phi-3-mini-4k-instruct on both the short-context (MMLU, GSM8K, HumanEval) and long-context (GovReport) tasks, while still having the 2048 window size of its SWA layer and maintaining the linear complexity for efficient processing of long documents. Details of data statistics and evaluation setup for long context tasks are included in Appendix F.
@@
Benchmark Llama-3 Mistral Mamba Mamba-SWA-MLP Mamba-MLP Samba
1.6B 1.6B 1.8B 1.6B 1.9B 1.7B
@@
To examine the different hybridization strategies mentioned in Section 2.1, we train 6 models with around 1.7B parameters on the Phi2 (Li et al., 2023) dataset with 230B tokens and evaluate them in the full suite of 15 downstream benchmarks to have a holistic assessment of hybrid and purebred architectures. As shown in Table 2, Samba demonstrates superior performance on a diverse set of tasks, including commonsense reasoning (ARC-Challenge), language understanding (MMLU, SQuAD), TruthfulQA and code generation (HumanEval, MBPP). It outperforms both the pure attention-based and SSM-based models in most tasks and achieves the best average performance. By comparing the performance of Mamba-MLP and Mamba in Table 2, we can observe that replacing Mamba blocks with MLPs does not harm common sense reasoning ability, but its performance in language understanding and complex reasoning ability, such as coding and mathematical reasoning, degenerates significantly. We can also see that pure Mamba models fall short on retrieval intensive tasks such as SQuAD due to their lack of precise memory retrieval ability. The best results are achieved through the combination of the attention and Mamba modules, as shown with our Samba architecture. We can also notice that Mamba-SWA-MLP has significantly better performance on GSM8K, potentially resulting from a closer collaboration between the Mamba and the SWA layers. The distinct downstream performances of different hybridization strategies pose interesting future work for developing task-adaptive dynamic architectures.
3.2 Exploration on Hybridizing Attention and Linear Recurrence